### v2

# Groups admitting Wirtinger presentations and Gromov hyperbolic groups

Abstract.
Twisted Wirtinger presentations are generalizations of the classical Wirtinger presentations of knot and link groups. In this paper, we prove that if a finitely generated group admitting a twisted Wirtinger presentation is Gromov hyperbolic, then its second rational homology group vanishes. Moreover, if the group is torsion-free, then its second integral homology group also vanishes.
Key words and phrases:
Wirtinger presentation, hyperbolic group, group homology
2020 Mathematics Subject Classification:
Primary 20F05,20F67,20J06; Secondary 20F65
1. Introduction
We say that a group presentation $\langle X\mid R\rangle$ is a twisted Wirtinger presentation if each relation in $R$ is of the form
$w^{-1}xw=y^{\epsilon}\quad(x,y\in X,w\in F(X),\epsilon\in\{\pm 1\}),$
where $F(X)$ is the free group on $X$. In particular, if all the relations in $R$ are of the form
$w^{-1}xw=y\quad(x,y\in X,w\in F(X)),$
then we say that $\langle X\mid R\rangle$ is a Wirtinger presentation. Both Wirtinger and twisted Wirtinger presentations generalize the classical Wirtinger presentations of knot and link groups. Groups admitting Wirtinger presentations are known by various names. Among others, they are called Wirtinger groups in [Yajima], labelled oriented graph groups (LOG groups in short) in [Howie], and C-groups in [Kulikov, MR1307063]. Groups admitting twisted Wirtinger presentations are referred to as non-orientable C-groups in [MR1392843], where a characterization of these groups in terms of group homology was established.
Examples of groups that admit Wirtinger presentations include free and free abelian groups, braid groups and pure braid groups, Artin groups, Thompson’s group $F$, and higher-dimensional knot groups. As for twisted Wirtinger presentations, examples include the infinite dihedral group, elementary abelian 22-groups, Coxeter groups, twisted Artin groups, and cactus groups.
More generally, let $M$ be a closed smooth $n$-dimensional manifold smoothly embedded in $\mathbb{R}^{n+2}$. Then the fundamental group of the complement $\mathbb{R}^{n+2}\setminus M$ admits a twisted Wirtinger presentation, and it admits a Wirtinger presentation if $M$ is orientable [MR1841759].
Gromov hyperbolic groups—also known as word hyperbolic or simply hyperbolic groups—were introduced and developed by Gromov [MR0919829], and are among the most fundamental objects in geometric group theory. Roughly speaking, a finitely generated group is Gromov hyperbolic if its Cayley graph behaves like a negatively curved space. For the precise definition and properties of Gromov hyperbolic groups, see [MR1744486, MR1086648] for instance.
Free groups $F_{n}$ are Gromov hyperbolic, whereas free abelian groups $\mathbb{Z}^{n}$ $(n\geq 2)$ are not. Necessary and sufficient conditions for Coxeter groups to be Gromov hyperbolic were obtained in  [MR2636665] (see also [MR2360474]). A large number of Coxeter groups are Gromov hyperbolic, while others are not. These facts imply that finitely generated groups admitting twisted Wirtinger presentations can be either Gromov hyperbolic or not, depending on the specific case.
The purpose of this paper is to provide a simple group homological obstruction to Gromov hyperbolicity for finitely generated groups that admit twisted Wirtinger presentations, which is a consequence of the following result:
Theorem 1.
Let $G$ be a group admitting a twisted Wirtinger presentation. If $G$ does not contain a subgroup isomorphic to $\mathbb{Z}\times\mathbb{Z}$, then the second rational homology group $H_{2}(G;\mathbb{Q})$ is trivial. Moreover, if the group is torsion-free, then the second integral homology group $H_{2}(G;\mathbb{Z})$ is also trivial.
It is well-known that a Gromov hyperbolic group does not contain a subgroup isomorphic to $\mathbb{Z}\times\mathbb{Z}$ (cf. [MR1744486, MR1086648]). The following corollary follows immediately from Theorem 1:
Corollary.
Let $G$ be a finitely generated group admitting a twisted Wirtinger presentation. If $G$ is Gromov hyperbolic, then the second rational homology group $H_{2}(G;\mathbb{Q})$ is trivial. Moreover, if $G$ is both Gromov hyperbolic and torsion-free, then the second integral homology group $H_{2}(G;\mathbb{Z})$ is also trivial.
We refer to [MR672956] for the standard facts about group homology. It is worth noting that any finitely generated abelian group AA can be realized as the second homology group $H_{2}(G;\mathbb{Z})$ of a group $G$ admitting a Wirtinger presentation [MR0635591]. Hence, a large number of groups admitting twisted Wirtinger presentations cannot be Gromov hyperbolic. Conversely, many Gromov hyperbolic groups have positive second Betti number, and such groups cannot admit twisted Wirtinger presentations.
2. Proof
[…]
$g\wedge h\coloneqq[g|h]-[h|g]\in H_{2}(G),$
where the right-hand side represents a normalized 22-cycle of $G$. This homology class is sometimes referred to as the Pontryagin product of $g,h\in G$. It is natural with respect to group homomorphisms. Namely, for a group homomorphism $f\colon G\to H$ and commuting elements $g,h\in G$, one has
$f(g)\wedge f(h)=f_{*}(g\wedge h)\in H_{2}(H).$
As a consequence of the naturality, for any commuting elements $g,h\in G$, consider the homomorphism $\phi\colon\mathbb{Z}\times\mathbb{Z}$ defined by $\phi(1,0)=g$, $\phi(0,1)=h$. Then the homology class $g\wedge h$ is exactly the image of $(1,0)\wedge(0,1)\in H_{2}(\mathbb{Z}\times\mathbb{Z})$ under the induced homomorphism $\phi_{*}\colon H_{2}(\mathbb{Z}\times\mathbb{Z})\to H_{2}(G)$, that is
($\star$) $g\wedge h=\phi_{*}((1,0)\wedge(0,1))\in H_{2}(G).$
Note that the integral homology class $(1,0)\wedge(0,1)\in H_{2}(\mathbb{Z}\times\mathbb{Z})$ can be shown to be a generator of $H_{2}(\mathbb{Z}\times\mathbb{Z})\cong\mathbb{Z}$. For the properties of the Pontryagin product, see [akita-even-Artin].
[…]
Let $G$ be a group admitting a twisted Wirtinger presentation. For any homology class $u\in H_{2}(G)$, there exist commuting elements $g,h\in G$ such that $u=g\wedge h\in H_{2}(G)$.
Proof of Theorem 1.
Let $G$ be a group. Suppose that
  1. (1)
    $G$ does not contain a subgroup isomorphic to $\mathbb{Z}\times\mathbb{Z}$,
  2. (2)
    $G$ admits a twisted Wirtinger presentation.
Let $\phi\colon\mathbb{Z}\times\mathbb{Z}\to G$ be a group homomorphism, and let $A\subset G$ be the image of $\phi$. It follows from the condition (1) that AA is an abelian group generated by two elements and of rank at most one. For such an abelian group AA, the second rational homology group $H_{2}(A;\mathbb{Q})=H_{2}(A)\otimes\mathbb{Q}$ is trivial. This implies that the induced homomorphism $\phi_{*}\colon H_{2}(\mathbb{Z}\times\mathbb{Z};\mathbb{Q})\to H_{2}(G;\mathbb{Q})$ is the zero map, since $\phi_{*}$ factors through $H_{2}(A;\mathbb{Q})$ as
$H_{2}(\mathbb{Z}\times\mathbb{Z};\mathbb{Q})\to H_{2}(A;\mathbb{Q})\to H_{2}(G;\mathbb{Q}).$
In view of the identity ($\star$2), the rational homology class $g\wedge h\in H_{2}(G;\mathbb{Q})$ vanishes for any commuting elements $g,h\in G$. On the other hand, by Theorem 2, every element of $H_{2}(G;\mathbb{Q})=H_{2}(G)\otimes\mathbb{Q}$ is a rational multiple of an element of the form $g\wedge h\in H_{2}(G;\mathbb{Q})$ for some commuting $g,h\in G$. Thus we conclude that $H_{2}(G;\mathbb{Q})=0$. This completes the proof of the first half of the theorem.
We now assume, in addition, that $G$ is torsion-free. Let AA be the image of a homomorphism $\phi\colon\mathbb{Z}\times\mathbb{Z}\to G$ as before. Then, by the condition (1), the subgroup AA must be either trivial or infinite cyclic. For such a group AA, the second integral homology $H_{2}(A)$ vanishes. This implies that the induced homomorphism $\phi_{*}\colon H_{2}(\mathbb{Z}\times\mathbb{Z})\to H_{2}(G)$ is the zero map. The second half of the theorem can be proved in a similar manner. ∎
Acknowledgement.
The author was partially supported by JSPS KAKENHI Grant Number 24K06727.